What is the volume in liters of 4.8 moles methane gas, CH 4 , at 30 ∘ C and 1.60 atm ? Express your answer using two significant figures.
Answer #1

: Given : T = 30+ 273.15 = 303.15 K    P = 1.60 atm

we know that R = 0.08206 Latm/mol-K

Using ideal gas equation PV = nRT we have

V= nRT/P = 4.8(mol) * 0.08206(Latmmol-1K-1) * 303.15K / 1.60atm

= 74.629467L

After rounding off to two significant figures: V = 75L